# Light single-gluon hybrid states with various (exotic) quantum numbers

### v1

@@ I Introduction
In this paper we shall systematically investigate the single-gluon hybrid states through the QCD sum rule method. Especially, we shall concentrate on the hybrid states with the exotic quantum number $J^{PC}=1^{-+}$ and update the previous calculations on their mass spectrum Latorre:1985tg , given that some QCD parameters have been significantly changed in recent years. We shall also update our previous calculations on their decay properties Chen:2010ic ; Huang:2010dc ; Chen:2022qpd . Assuming their quark-gluon contents to be either $\bar{q}qg$ ($q=u/d$) or $\bar{s}sg$ and their isospin to be either $I=1$ or $I=0$, we shall calculate their masses and widths to be:
@@
This paper is organized as follows. In Sec. II we construct twenty-four single-gluon hybrid currents with various (exotic) quantum numbers, eighteen of which are used to perform QCD sum rule analyses in Sec. III. After performing numerical analyses in Sec. IV, we calculate the masses of forty-four single-gluon hybrid states with the quark-gluon contents $\bar{q}qg$ ($q=u/d$) and $\bar{s}sg$. The obtained results are summarized and discussed in Sec. V. Based on these results, we update our previous calculations on the decay properties of the $J^{PC}=1^{-+}$ hybrid states Chen:2010ic ; Huang:2010dc ; Chen:2022qpd , and the obtained results are also summarized in Sec. V.
@@ II Single-gluon hybrid currents
Especially, we shall concentrate on the fifth current $J^{\mu}_{1^{-+}}$ with the exotic quantum number $J^{PC}=1^{-+}$. In the above expressions $\{\alpha\beta\}/\{\alpha_{1}\beta_{1}\}/\{\alpha_{2}\beta_{2}\}$ are antisymmetric Lorentz pairs. The four currents $J^{\mu,\alpha\beta}_{2^{-+}}$, $\widetilde{J}^{\mu,\alpha\beta}_{2^{++}}$, $J^{\mu,\alpha\beta}_{2^{+-}}$, and $\widetilde{J}^{\mu,\alpha\beta}_{2^{--}}$ all contain three Lorentz indices with the mixed symmetry, so their spin-2 components can not be easily extracted. We shall not investigate these four currents in the present study, and we refer to Ref. Wang:2023whb for detailed discussions. The symbol $\mathcal{A}[\cdots]$ represents anti-symmetrization in the set $\{\alpha\beta\}$, which can be done by multiplying the projection operator
@@ III QCD sum rule analyses
IV Numerical Analyses
In this section we use the OPE spectral densities given in Eq. (III) as well as those listed in Appendix A to perform numerical analyses. We use the spectral density $\rho^{\bar{s}sg}_{1^{-+}}(s)$ given in Eq. (III) as an example. It is extracted from the current $J^{\mu}_{1^{-+}}$ with the quark-gluon content $\bar{s}sg$, which couples to the state
@@ V Summary and Discussions
Figure 5: Decay mechanisms of the single-gluon hybrid states through (a) the normal process with one quark-antiquark pair excited from the valence gluon, and (b,c) the abnormal processes with the $\eta^{(\prime)}$ produced by the QCD axial anomaly.
@@
$f_{|\bar{s}sg;0^{+}1^{-+}\rangle}$ $=$ $0.300^{+0.063}_{-0.058}{\rm~{}GeV}\,.$
As shown in Fig. 5(a), a single-gluon hybrid state can decay after exciting one $\bar{q}q/\bar{s}s$ pair from the valence gluon, followed by reorganizing two color-octet $\bar{q}q/\bar{s}s$ pairs into two color-singlet mesons. This decay process has been systematically studied in Refs. Chen:2010ic ; Huang:2010dc for the $J^{PC}=1^{-+}$ hybrid states through the QCD sum rule method, and in the present study we update these calculations. Besides the “normal” decay process depicted in Fig. 5(a), there also exist the “abnormal” decay processes depicted in Figs. 5(b,c), where the $\eta^{(\prime)}$ mesons are produced by the QCD axial anomaly. These abnormal decay processes have been systematically studied in Ref. Chen:2022qpd , and in the present study we also update these calculations. The obtained results are summarized in Table 3, and especially,

### v2

@@ I Introduction
Besides the exotic quantum numbers $J^{PC}=0^{+-}/1^{-+}/2^{+-}$, the light single-gluon hybrid states with other quantum numbers have not been well studied in the literature. Accordingly, in this paper we shall systematically investigate the single-gluon hybrid states with various (exotic) quantum numbers through the QCD sum rule method. Especially, we shall concentrate on the hybrid states with the exotic quantum number $J^{PC}=1^{-+}$ and update the previous calculations on their mass spectrum Latorre:1985tg , given that some QCD parameters have been significantly changed in recent years. We shall also update our previous calculations on their decay properties Chen:2010ic ; Huang:2010dc ; Chen:2022qpd , with more decay channels taken into account (see the caption of Table 3). Assuming their quark-gluon contents to be either $\bar{q}qg$ ($q=u/d$) or $\bar{s}sg$ and their isospin to be either $I=1$ or $I=0$, we shall calculate their masses and widths to be:
@@
This paper is organized as follows. In Sec. II we construct twenty-four single-gluon hybrid currents with various (exotic) quantum numbers. In Sec. III we use eighteen of them to perform QCD sum rule analyses, and calculate masses of forty-four single-gluon hybrid states with the quark-gluon contents $\bar{q}qg$ ($q=u/d$) and $\bar{s}sg$. Based on these results, we systematically study the decay properties of the $J^{PC}=1^{-+}$ hybrid states in Sec. IV. The obtained results are summarized in Sec. V.
@@ II Single-gluon hybrid currents
Especially, we shall concentrate on the fifth current $J^{\mu}_{1^{-+}}$ with the exotic quantum number $J^{PC}=1^{-+}$, while this current also contains the $J^{PC}=0^{++}$ component, so we need to separate the $J^{PC}=1^{-+}$ and $0^{++}$ components in the calculations, as discussed below. In the above expressions $\{\alpha\beta\}/\{\alpha_{1}\beta_{1}\}/\{\alpha_{2}\beta_{2}\}$ are antisymmetric Lorentz pairs. The four currents $J^{\mu,\alpha\beta}_{2^{-+}}$, $\widetilde{J}^{\mu,\alpha\beta}_{2^{++}}$, $J^{\mu,\alpha\beta}_{2^{+-}}$, and $\widetilde{J}^{\mu,\alpha\beta}_{2^{--}}$ all contain three Lorentz indices with the mixed symmetry, so their spin-2 components can not be easily extracted. We shall not investigate these four currents in the present study, and we refer to Ref. Wang:2023whb for detailed discussions. The symbol $\mathcal{A}[\cdots]$ represents anti-symmetrization in the set $\{\alpha\beta\}$, which can be done by multiplying the projection operator
@@ III QCD sum rule analyses
We use the spectral density $\rho^{\bar{s}sg}_{1^{-+}}(s)$ given in Eq. (III) to perform numerical analyses. It is extracted from the current $J^{\mu}_{1^{-+}}$ with the quark-gluon content $\bar{s}sg$, which couples to the state
@@
IV Decay properties of the $J^{PC}=1^{-+}$ hybrid states
In this section we systematically study the decay properties of the $J^{PC}=1^{-+}$ hybrid states, whose masses and decay constants have been calculated in the previous section. Since we do not differentiate the up and down quarks within the QCD sum rule method, the masses and decay constants of the states in the same isospin multiplet are calculated to be the same:
$f_{|\bar{q}qg;1^{-}1^{-+}\rangle}=f_{|\bar{q}qg;0^{+}1^{-+}\rangle}$ $=$ $0.243^{+0.057}_{-0.052}{\rm~{}GeV}^{4}\,,$
$M_{|\bar{s}sg;0^{+}1^{-+}\rangle}$ $=$ $1.84^{+0.14}_{-0.15}{\rm~{}GeV}\,,$
$f_{|\bar{s}sg;0^{+}1^{-+}\rangle}$ $=$ $0.300^{+0.063}_{-0.058}{\rm~{}GeV}^{4}\,.$
As shown in Fig. 5(a), a single-gluon hybrid state can decay after exciting one $\bar{q}q/\bar{s}s$ pair from the valence gluon, followed by reorganizing two color-octet $\bar{q}q/\bar{s}s$ pairs into two color-singlet mesons. This decay process has been systematically studied in Refs. Chen:2010ic ; Huang:2010dc for the $J^{PC}=1^{-+}$ hybrid states through the QCD sum rule method, and in the present study we update these calculations. Besides the “normal” decay process depicted in Fig. 5(a), there also exist the “abnormal” decay processes depicted in Figs. 5(b,c), where the $\eta^{(\prime)}$ mesons are produced by the QCD axial anomaly. These abnormal decay processes have been systematically studied in Ref. Chen:2022qpd , and in the present study we also update these calculations.
@@ IV Decay properties of the $J^{PC}=1^{-+}$ hybrid states
Figure 5: Decay mechanisms of the single-gluon hybrid states through (a) the normal process with one quark-antiquark pair excited from the valence gluon, and (b,c) the abnormal processes with the $\eta/\eta^{\prime}$ produced by the QCD axial anomaly.
[…]
IV.1 Normal decay process
In this subsection we use the decay mode
$\pi_{1}\equiv|\bar{q}qg;1^{-}1^{-+}\rangle\rightarrow\rho\pi\,,$ (57)
as an example to study the normal decay process depicted in Fig. 5(a) through the three-point correlation function
$T_{\mu\nu}(p,k,q)$ $=$ $\int d^{4}xd^{4}ye^{ikx}e^{iqy}\times$
$~{}~{}~{}~{}~{}~{}\langle 0|{\mathbb{T}}[J_{\nu}^{\rho^{-}}(x)J_{5}^{\pi^{+}}(y)J^{\mu\dagger}_{1^{-+}}(0)]|0\rangle\,,$
where $p$, $k$ and $q$ are the momenta of $\pi_{1}\equiv|\bar{q}qg;1^{-}1^{-+}\rangle$, $\rho^{-}$, and $\pi^{+}$, respectively. The current $J^{\mu}_{1^{-+}}$ has been defined in Eq. (11), and we select the isovector neutral-charged one
$J^{\mu}_{1^{-+}}\rightarrow{1\over\sqrt{2}}\left(\bar{u}_{a}\lambda_{n}^{ab}\gamma_{\beta}u_{b}-\bar{d}_{a}\lambda_{n}^{ab}\gamma_{\beta}d_{b}\right)g_{s}G_{n}^{\mu\beta}\,.$ (59)
The negative-charged vector current $J_{\mu}^{\rho^{-}}=\bar{u}\gamma_{\mu}d$ and the positive-charged pseudoscalar current $J_{5}^{\pi^{+}}=\bar{d}\gamma_{5}u$ respectively couple to the vector meson $\rho^{-}$ and the pseudoscalar meson $\pi^{+}$ through
$\langle 0|J_{\mu}^{\rho^{-}}|\rho^{-}(k,\epsilon)\rangle$ $=$ $m_{\rho}f_{\rho}\epsilon_{\mu}\,,$ (60)
$\langle 0|J_{5}^{\pi^{+}}|\pi(q)\rangle$ $=$ $f_{\pi}^{\prime}={2i\langle\bar{q}q\rangle\over f_{\pi}}\,,$ (61)
with Zhu:1999wg ; Reinders:1984sr ; Zhu:1998bm :
$m_{\pi}={140\rm MeV}\,,~{}~{}~{}f_{\pi}=131{\rm MeV}\,,$ (62)
$m_{\rho}=770{\rm MeV}\,,~{}~{}~{}f_{\rho}=220{\rm MeV}\,.$
At the phenomenological side we write $T_{\mu\nu}(p,k,q)$ as
$T_{\mu\nu}(p,k,q)$ $=$ $g_{\rho\pi}\epsilon_{\mu\nu\alpha\beta}{q^{\alpha}k^{\beta}}$
$\times$ ${f_{\pi_{1}}f_{\rho}m_{\rho}f_{\pi}^{\prime}\over(m_{\pi_{1}}^{2}-p^{2})(m_{\rho}^{2}-k^{2})(m_{\pi}^{2}-q^{2})}\,,$
where the coupling constant $g_{\rho\pi}$ is defined through the Lagrangian
${\cal L}=g_{\rho\pi}\epsilon_{\mu\nu\alpha\beta}~{}\pi_{1}^{0\mu}~{}\partial^{\alpha}\pi^{+}~{}\partial^{\beta}\rho^{-\nu}+\cdots\,.$ (64)
At the QCD side we calculate $T_{\mu\nu}(p,k,q)$ using the method of operator product expansion. We work at the pion pole and choose the terms divergent at the $q^{2}\rightarrow 0$ limit to derive
$T_{\mu\nu}(p,k,q)={\epsilon_{\mu\nu\alpha\beta}{q^{\alpha}k^{\beta}}\over q^{2}}\times$ (65)
$~{}~{}~{}\Big{(}{\langle g_{s}\bar{q}\sigma Gq\rangle\over 6\sqrt{2}}\big{(}{3\over p^{2}}+{1\over k^{2}}\big{)}-{\langle\bar{q}q\rangle\langle g_{s}^{2}GG\rangle\over 18\sqrt{2}}\big{(}{1\over p^{4}}+{1\over k^{4}}\big{)}\Big{)}\,.$
We compare Eq. (IV.1) and Eq. (65) to calculate the coupling constant $g_{\rho\pi}$. After setting $p^{2}=k^{2}$ and performing the Borel transformation once $\mathcal{B}(p^{2}=k^{2}\rightarrow T^{2})$, we arrive at
$-g_{\rho\pi}{f_{\pi_{1}}f_{\rho}m_{\rho}f_{\pi}^{\prime}\over m_{\rho}^{2}-m_{\pi_{1}}^{2}}\Big{(}e^{-m_{\pi_{1}}^{2}/T^{2}}-e^{-m_{\rho}^{2}/T^{2}}\Big{)}$ (66)
$=$ $-{2\langle g\bar{q}\sigma Gq\rangle\over 3\sqrt{2}}-{\langle\bar{q}q\rangle\langle g_{s}^{2}GG\rangle\over 9\sqrt{2}}{1\over T^{2}}\,.$
The formula of the decay width reads
$\Gamma(\pi_{1}^{0}\rightarrow\rho^{+}\pi^{-}+\rho^{-}\pi^{+})=2\times{g_{\rho\pi}^{2}\over 12\pi}|\vec{q}_{\pi}|^{3}\,,$ (67)
where $\vec{q}_{\pi}$ is the three-momentum of the final state $\pi$. Numerically, we obtain
$g_{\rho\pi}$ $=$ ${4.08^{+2.40}_{-1.83}~{}\rm GeV^{-1}}\,,$ (68)
$\Gamma(\pi_{1}\rightarrow\rho\pi)$ $=$ ${242^{+310}_{-179}~{}\rm MeV}\,.$ (69)
IV.2 Abnormal decay process
In this subsection we use the decay mode
[…]
as an example to study the abnormal decay process depicted in Fig. 5(c), with $\eta^{\prime}$ produced by the QCD $U(1)_{A}$ anomaly. Before doing this, let us shortly introduce the two-angle mixing formalism to describe the $\eta$ and $\eta^{\prime}$ mesons Leutwyler:1997yr ; Kaiser:1998ds ; Escribano:2005qq ; Escribano:2015nra ; Escribano:2015yup ; Schechter:1992iz ; Kiselev:1992ms ; Herrera-Siklody:1997pgy ; Bass:2018xmz ; Bali:2021qem :
$|\eta\rangle$ $=$ $\cos\theta_{8}|\eta_{8}\rangle-\sin\theta_{0}|\eta_{0}\rangle+\cdots\,,$ (71)
$|\eta^{\prime}\rangle$ $=$ $\sin\theta_{8}|\eta_{8}\rangle+\cos\theta_{0}|\eta_{0}\rangle+\cdots\,,$
with
$|\eta_{8}\rangle$ $=$ $|u\bar{u}+d\bar{d}-2s\bar{s}\rangle/\sqrt{6}\,,$ (72)
$|\eta_{0}\rangle$ $=$ $|u\bar{u}+d\bar{d}+s\bar{s}\rangle/\sqrt{3}\,,$
and $\cdots$ are contributions from the pseudoscalar glueballs and charmonium, etc.
The octet and singlet axial-vector currents are defined as
$A_{\mu}^{8}$ $=$ $\left({\bar{u}\gamma_{\mu}\gamma_{5}u+\bar{d}\gamma_{\mu}\gamma_{5}d-2\bar{s}\gamma_{\mu}\gamma_{5}s}\right)/\sqrt{12}\,,$ (73)
$A_{\mu}^{0}$ $=$ $\left({\bar{u}\gamma_{\mu}\gamma_{5}u+\bar{d}\gamma_{\mu}\gamma_{5}d+\bar{s}\gamma_{\mu}\gamma_{5}s}\right)/\sqrt{6}\,.$
These two currents couple to the $\eta$ and $\eta^{\prime}$ mesons through
$\langle 0|A_{\mu}^{a}|P(k)\rangle=ik_{\mu}f_{P}^{a}\,,$ (74)
where $f_{P}^{a}$ ($a=8,0;\,P=\eta,\eta^{\prime}$) is the matrix for the decay constants, defined as
$\left(\begin{array}[]{cc}f_{\eta}^{8}&f_{\eta}^{0}\\ f_{\eta^{\prime}}^{8}&f_{\eta^{\prime}}^{0}\end{array}\right)=\left(\begin{array}[]{cc}f_{8}\cos\theta_{8}&-f_{0}\sin\theta_{0}\\ f_{8}\sin\theta_{8}&f_{0}\cos\theta_{0}\end{array}\right)\,.$ (75)
To simply our notations, we further construct the axial-vector currents
$J_{\mu}^{\eta}=A_{\mu}^{8}+t_{\eta}A_{\mu}^{0}\,,$ (76)
$J_{\mu}^{\eta\prime}=A_{\mu}^{8}+t_{\eta^{\prime}}A_{\mu}^{0}\,,$ (77)
which couple to the $\eta$ and $\eta^{\prime}$ mesons through
$\langle 0|J_{\mu}^{\eta}|\eta(k)\rangle$ $=$ $ik_{\mu}g_{\eta}\,,$
$\langle 0|J_{\mu}^{\eta\prime}|\eta^{\prime}(k)\rangle$ $=$ $ik_{\mu}g_{\eta^{\prime}}\,,$ (78)
$\langle 0|J_{\mu}^{\eta}|\eta^{\prime}(k)\rangle$ $=$ $\langle 0|J_{\mu}^{\eta\prime}|\eta(k)\rangle=0\,,$
with
$g_{\eta}$ $=$ $f_{\eta}^{8}-f_{\eta}^{0}f_{\eta^{\prime}}^{8}/f_{\eta^{\prime}}^{0}\,,$
$g_{\eta^{\prime}}$ $=$ $f_{\eta^{\prime}}^{8}-f_{\eta^{\prime}}^{0}f_{\eta}^{8}/f_{\eta}^{0}\,,$ (79)
$t_{\eta}$ $=$ $-f_{\eta^{\prime}}^{8}/f_{\eta^{\prime}}^{0}\,,$
$t_{\eta^{\prime}}$ $=$ $-f_{\eta}^{8}/f_{\eta}^{0}\,.$
We shall use the following values in the calculations Ali:1997ex ; Feldmann:1997vc :
$\theta_{8}$ $=$ $-22.2^{\circ}\,,$
$\theta_{0}$ $=$ $-9.1^{\circ}\,,$
$f_{8}$ $=$ $168{\rm~{}MeV}\,,$ (80)
$f_{0}$ $=$ $157{\rm~{}MeV}\,.$
The conversion of the gluons into the $\eta$ and $\eta^{\prime}$ mesons can be described through the QCD $U(1)_{A}$ anomaly as Voloshin:1980zf ; Akhoury:1987ed ; Castoldi:1988dm ; Chao:1989yp ; Ball:1995zv ; Ali:1997ex :
$\langle 0|{\alpha_{s}\over 4\pi}G^{\alpha\beta}_{n}\tilde{G}_{n,\alpha\beta}|\eta\rangle$ $=$ $m^{2}_{\eta}f_{\eta}\,,$ (81)
$\langle 0|{\alpha_{s}\over 4\pi}G^{\alpha\beta}_{n}\tilde{G}_{n,\alpha\beta}|\eta^{\prime}\rangle$ $=$ $m^{2}_{\eta^{\prime}}f_{\eta^{\prime}}\,,$ (82)
with
$f_{\eta}$ $=$ ${f_{8}\over\sqrt{6}}\cos\theta_{8}-{f_{0}\over\sqrt{3}}\sin\theta_{0}\,,$ (83)
$f_{\eta^{\prime}}$ $=$ ${f_{8}\over\sqrt{6}}\sin\theta_{8}+{f_{0}\over\sqrt{3}}\cos\theta_{0}\,.$ (84)
[…]
$T^{\prime}_{\mu\nu}(p,k,q)=\int d^{4}xe^{-ikx}\langle 0|{\mathbb{T}}[J^{\mu}_{1^{-+}}(0)J_{\nu}^{\eta\dagger}(x)]|\eta^{\prime}\rangle\,,$ (85)
where $p$, $k$, and $q$ are the momenta of $\eta_{1}\equiv|\bar{s}sg;0^{+}1^{-+}\rangle$, $\eta$, and $\eta^{\prime}$, respectively. The current $J_{\nu}^{\eta}$ has been defined in Eq. (76). The current $J^{\mu}_{1^{-+}}$ has been defined in Eq. (11), and we set its quark content to be
$J^{\mu}_{1^{-+}}\rightarrow\bar{s}_{a}\lambda_{n}^{ab}\gamma_{\beta}s_{b}g_{s}G_{n}^{\mu\beta}\,.$ (86)
At the phenomenological side we write $T^{\prime}_{\mu\nu}(p,k,q)$ as
$T^{\prime}_{\mu\nu}(p,k,q)=g_{\eta\eta^{\prime}}k_{\mu}k_{\nu}~{}{f_{\eta_{1}}g_{\eta}\over(m_{\eta_{1}}^{2}-p^{2})(m_{\eta}^{2}-k^{2})}\,,$ (87)
where the coupling constant $g_{\eta\pi}$ is defined through the Lagrangian
$\mathcal{L}^{\prime}=g_{\eta\eta^{\prime}}~{}\eta_{1}^{\mu}~{}(\partial_{\mu}\eta)~{}\eta^{\prime}\,.$ (88)
At the QCD side we calculate $T^{\prime}_{\mu\nu}(p,k,q)$ using the method of operator product expansion to be
$T^{\prime}_{\mu\nu}(p,k,q)$
$=$ $\theta_{s}k_{\mu}k_{\nu}~{}\Big{(}-{2m^{2}_{\eta^{\prime}}f_{\eta^{\prime}}\over 3k^{2}}-{4\pi^{2}m^{2}_{\eta^{\prime}}f_{\eta^{\prime}}m_{s}\langle\bar{s}s\rangle\over 3k^{6}}\Big{)}\,,$
where $\theta_{s}=-{1/\sqrt{3}}+{t_{\eta}/\sqrt{6}}$ describes the $s\bar{s}$ component contained in the current $J_{\nu}^{\eta}$.
After setting $p^{2}=k^{2}$ and performing the Borel transformation once $\mathcal{B}(p^{2}=k^{2}\rightarrow T^{2})$, we arrive at
$g_{\eta\eta^{\prime}}{f_{\eta_{1}}g_{\eta}\over m_{\eta_{1}}^{2}-m_{\eta}^{2}}\left(e^{-m_{\eta}^{2}/M_{B}^{2}}-e^{-m_{\eta_{1}}^{2}/M_{B}^{2}}\right)$ (90)
$=$ ${2\theta_{s}m^{2}_{\eta^{\prime}}f_{\eta^{\prime}}\over 3}+{2\pi^{2}\theta_{s}m^{2}_{\eta^{\prime}}f_{\eta^{\prime}}m_{s}\langle\bar{s}s\rangle\over 3}{1\over M_{B}^{4}}\,.$
The formula of the decay width reads
$\Gamma(\eta_{1}\rightarrow\eta\eta^{\prime})={g_{\eta\eta^{\prime}}^{2}\over 24\pi m_{\eta_{1}}^{2}}|\vec{q}_{\eta}|^{3}\,.$ (91)
where $\vec{q}_{\eta}$ is the three-momentum of the final state $\eta$. Numerically, we obtain
$g_{\eta\eta^{\prime}}$ $=$ $3.08^{+1.30}_{-0.91}\ \,,$ (92)
[…]
Similarly, we study the other normal and abnormal decay processes. The obtained results are summarized in Table 3.
@@ V Summary and Discussions
$f_{|\bar{s}sg;0^{+}1^{-+}\rangle}$ $=$ $0.300^{+0.063}_{-0.058}{\rm~{}GeV}^{4}\,.$
There have been a lot of Lattice QCD calculations on the $I^{G}J^{PC}=1^{-}1^{-+}$ hybrid state in the past fifty years Griffiths:1983ah ; Michael:1985ne ; Perantonis:1990dy ; Foster:1998wu ; Juge:2002br ; Bali:2003jq , and especially, the Hadron Spectrum collaboration have performed exhaustive analyses Dudek:2009qf ; Dudek:2010wm ; Dudek:2011bn ; HadronSpectrum:2012gic ; Dudek:2013yja ; Woss:2020ayi ; Briceno:2017max . C. A. Meyer and E. S. Swanson summarized these results, and the naive extrapolation of its mass, calculated by Lattice QCD, to the physical pion mass turns out to be approximately 1.6 GeV Meyer:2015eta . Therefore, our QCD sum rule calculation is well consistent with the Lattice QCD calculations.
Based on the mass calculations, we systematically study the decay properties of the $J^{PC}=1^{-+}$ hybrid states. We have considered the normal decay process depicted in Fig. 5(a). We have also considered the abnormal decay processes depicted in Figs. 5(b,c), with the $\eta^{(\prime)}$ mesons produced by the QCD axial anomaly. The obtained results are summarized in Table 3, and especially,